# Nontrivial multi-product commutation relation for reducing $T$-count in sequential Pauli-based computation

### v1

@@ V Conclusion and outlook
Through numerical experiments, we demonstrate that circuits transformed using MCR-based unoptimization are challenging for existing $T$-count optimization compilers to optimize, suggesting that MCR offers circuit rewrites that remain unexplored by existing compilation strategies. We also compare the optimization performance results with and without MCR-based gate swapping. The comparison shows that while MCR-based gate swapping introduces additional circuit redundancy, it can occasionally trigger effective local optimization in small regions of the circuit. However, such cases become infrequent as the number of qubits grows, due to the exponential increase in the number of valid MCR candidates.
MCR involves circuit transformations that can increase or decrease the number of $T$ layers, thereby suggesting a broader class of rewrite rules beyond the reach of existing compiler strategies. This insight opens opportunities for compiler design, circuit synthesis, and algorithm-based optimization. A key direction for future work is to design a compiler that incorporates an MCR-based gate swapping technique toward further $T$-count reduction. However, simply adding such a rule to existing optimization passes is not sufficient, since the number of possible MCR candidates grows exponentially with the system size, making brute-force exploration computationally infeasible. Addressing this challenge requires developing a search strategy that partitions the circuit into manageable subcircuits and selectively explores promising regions. Once developed, the performance of such a compiler can be evaluated on representative datasets, such as the Reversible Benchmarks [39].

### v2

@@ V Conclusion and outlook
Through numerical experiments, we demonstrate that circuits transformed using MCR-based unoptimization are challenging for existing $T$-count optimization compilers to optimize, while our MCR-aware compiler can successfully eliminate the redundancies. We also compare the optimization performance results with and without MCR-based gate swapping. The comparison shows that while MCR-based gate swapping introduces additional circuit redundancy, it can occasionally trigger effective local optimization in small regions of the circuit. However, such cases become infrequent as the number of qubits grows, due to the exponential increase in the number of valid MCR candidates.
MCR involves circuit transformations that can increase or decrease the number of $T$ layers, thereby suggesting a broader class of rewrite rules beyond the reach of existing compiler strategies. This insight opens opportunities for compiler design, circuit synthesis, and algorithm-based optimization. A key direction for future work is to design a scalable compiler that more effectively exploits MCR-based gate swapping techniques toward further $T$-count reduction. However, simply adding such a rule to existing optimization passes is not sufficient, since the number of possible MCR candidates grows exponentially with the system size, making brute-force exploration computationally infeasible. Addressing this challenge requires developing a search strategy that partitions the circuit into manageable subcircuits and selectively explores promising regions. Once developed, the performance of such a compiler can be evaluated on representative datasets, such as the Reversible Benchmarks [40].